A gas is initially  at volume 40.0 L, pressure 600 mmHg and temperature 35 °C. Calculate the final pressure of the gas in atm if the volume and temperature are changed to 2500 mL and 225 K, respectively.

1
Expert's answer
2021-11-23T11:08:02-0500

P1=600/760=0.789atm P2=?

V1=40L V2=2.5L

T1=308K T2=225K


Using general gas law

P1V1/T1=P2V2/T2

P2=P1V1T2/V2T1

P2=(0.789×40×225)/2.5×308

P2=9.22atm.